OpenStudy (anonymous):

ameila sets her cd player to randomly select songs from the 4 cd's she loaded. disc a has 11 songs, disc b has 10 songs, disc c has 9 and dhad 14 songs. what is the probability that the first song she hears is from disk c?

OpenStudy (anonymous):

\[\frac{9}{11+10+9+14} = \frac{9}{34}\]

OpenStudy (anonymous):

9/34

OpenStudy (anonymous):

9/44

OpenStudy (anonymous):

9/44

OpenStudy (anonymous):

i mean

OpenStudy (anonymous):

lols yup i added wrong

OpenStudy (anonymous):

lol thanks